OpenStudy (anonymous):

So I just need confirmation that I'm doing this pre-calc right: Verify the identity. cotx minus pi divided by two. = -tan x

OpenStudy (anonymous):

\[\csc (x-\pi/2)=-tanx\]

OpenStudy (anonymous):

sorry cot not csc

OpenStudy (anonymous):

So I think the proper way to start it is with the even/odd identity of cot(-x)=-cotx but im not sure if the proper way to state this would be -cotxpi/2=-tanx or -cot(x+pi/2)

OpenStudy (anonymous):

The first thing I would do is simplify everything into terms of sin and cos 1/cot(x+pi/2) = -sinx/cos x

OpenStudy (anonymous):

Rather, cos(x+pi/2)/sin(x+pi/2)

OpenStudy (anonymous):

Wait, you've switched from a minus in the initial posting to a plus. Which is it?

OpenStudy (anonymous):

well it is originally a - but my teacher said to start with even/odd identities and im trying to figure out if that would automatically make it a positive

OpenStudy (anonymous):

For right now, I would leave it as stated, unless you foresee that being the first step in solving the problem.

OpenStudy (xapproachesinfinity):

cotx is odd so cot(-x)=-cotx

OpenStudy (anonymous):

but what about the pi/2?

OpenStudy (xapproachesinfinity):

cot(x-pi/2) is this the initial LHD

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Tangent and Cotangent are co-functions. \(\Large\rm \cot(90-x)=\tan x\) This is almost something that might help you :)

OpenStudy (xapproachesinfinity):

you don't need odd even identities you need cofunction identities or whatever they are called

OpenStudy (anonymous):

yeah but the part inside the parenthesis is backward from the confunction identity listed above

OpenStudy (xapproachesinfinity):

yeah like what @zepdrix said

OpenStudy (xapproachesinfinity):

oh my bad you need those odd even identities cot(x-pi/2)=-cot(pi/2-x)

OpenStudy (xapproachesinfinity):

so what's next

OpenStudy (anonymous):

divide by negative one would b the most logical, right?

OpenStudy (anonymous):

both sides of course

OpenStudy (xapproachesinfinity):

well look at what @zepdrix wrote above? doesn't that look the same

OpenStudy (anonymous):

Well in the rewritten side, what you wrote had a - in front of it and the identity does not. and so does tanx on the RHS so it would work

OpenStudy (xapproachesinfinity):

well you are trying to get -tanx we have cot(pi/2-x)=tanx so??

OpenStudy (xapproachesinfinity):

that negative is needed no?

OpenStudy (anonymous):

Yes I think so. Thanks guys!

OpenStudy (xapproachesinfinity):

good catch! you are welcome^_^